When styrene C8H8 burns in oxygen to form carbon dioxide and liquid water under standard-state conditions at 25°C, 42.62 kJ are released per gram of styrene. Find the standard enthalpy of formation of styrene at 25°C. (Given: ΔH°f[CO2(g)] = –393.5 kJ/mol, ΔH°f[H2O(l)] = –285.8 kJ/mol)
Answer #1

Answer:-

Given:-

standard enthalpy of formation of CO2(g) i.e \DeltaH0f [CO2(g)) ] = - 393.5 kJ/mol

standard enthalpy of formation of H2O(l) i.e \DeltaH0f [H2O(l) ] = - 285.8 kJ/mol

heat energy of 1 g styrene (C8H8) = - 42.62 kJ ( since heat energy released i.e exothermic reaction)

standard enthalpy of formation of styrene (C8H8) i.e \DeltaH0f [C8H8 ] = ?

As we know that

molar mass of styrene (C8H8) = 8 \times molar mass of C + 8 \times​​​​​​​ molar mass of H

molar mass of styrene (C8H8) = 8 \times​​​​​​​ 12.0107 + 8 \times​​​​​​​ 1.008

molar mass of styrene (C8H8) = 96.0856‬  + 8.064

molar mass of styrene (C8H8) = 104.15 g / mol

Also we know

1 mole styrene (C8H8) = 104.15 g

So

heat energy of 1 g styrene (C8H8) = -42.62 kJ

then

heat energy of 104.15 g styrene (C8H8) = 104.15 \times - 42.62 kJ

heat energy of 104.15 g styrene (C8H8) = - 4438.9 kJ

therefore when 1 mole  styrene (C8H8) burns in oxygen to form carbon dioxide and liquid water to produce - 4438.9 kJ heat energy which is standard heat of reaction i.e \DeltaH0rxn

So

standard heat of reaction (\DeltaH0rxn ) =   - 4438.9 kJ / mol

Since we know that

C8H8 + 10O2(g)  \rightarrow  8CO2(g) + 4H2O(l)

So

standard heat of reaction (\DeltaH0rxn ) = standard heat of products (\DeltaH0products ) - standard heat of reactants (\DeltaH0reactants )

standard heat of reaction (\DeltaH0rxn ) = (8 \times\DeltaH0f [CO2(g)) ] + 4 \times\DeltaH0f [H2O(l) ]   ) - ( \Delta​​​​​​​H0f [C8H8 ] )

- 4438.9 kJ / mol = (8 \times - 393.5 kJ/mol + 4 \times- 285.8 kJ/mol]   ) - ( \Delta​​​​​​​H0f [C8H8 ] )

- 4438.9 kJ / mol = ( - 3148.0 kJ/mol + - 1143.2 kJ/mol]   ) - ( \Delta​​​​​​​H0f [C8H8 ] )

- 4438.9 kJ / mol +  ( \Delta​​​​​​​H0f [C8H8 ] ) = ( - 3148.0 kJ/mol + - 1143.2 kJ/mol] )

- 4438.9 kJ / mol + \Delta​​​​​​​H0f [C8H8 ] = - 4291.2 kJ/mol]

\Delta​​​​​​​H0f [C8H8 ] = - 4291.2 kJ/mol] + 4438.9 kJ / mol

standard enthalpy of formation of styrene (C8H8) i.e \Delta​​​​​​​H0f [C8H8 ] = 147.7 kJ / mol (i.e the answer)